OpenStudy (anonymous):

please help. The midpoint of Uv is (5-11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

Directrix (directrix):

The midpoint coordinates are the averages of the x- and y-coordinates of the endpoints. So, solve these two equations to get the missing endpoint: (x+3)/2 = 5 (y+5)/2 = -11 @2cor57

OpenStudy (anonymous):

How do i solve them?

OpenStudy (anonymous):

@directrix

Directrix (directrix):

(x+3)/2 = 5 --> multiply both sides of the equation by 2 (x + 3) = 10 --> solve for x On this one, follow the same procedure: (y+5)/2 = -11 y = ? @2cor57